# From graph to Riesz continuity

### v2

@@ Abstract
We show that every graph continuous family of unbounded operators in a Hilbert space becomes Riesz continuous after multiplication by an appropriate family of unitary operators. This result provides a simple definition of the index for graph continuous families of Fredholm operators and leads to two corollaries for operators with compact resolvents: (1) the identity map between the spaces of such operators equipped with the Riesz and the graph topology is a homotopy equivalence; (2) every graph continuous family of such operators acting between fibers of Hilbert bundles becomes Riesz continuous in appropriate trivializations of the bundles.
For self-adjoint operators, multiplication by unitaries should be replaced by conjugation. In general, a graph continuous family of self-adjoint operators with compact resolvents cannot be made Riesz continuous by an appropriate conjugation. We obtain a partial analogue of the “trivialization” result above for self-adjoint operators and describe obstructions to existence of such a trivialization arising in the general case. This part of results is closely related to the recent work [$I_{2}$] of N. Ivanov.
footnotetext: Department of Mathematics, Technion – Israel Institute of Technology.footnotetext: This work was partially supported by ISF grants no. 431/20 and 844/19.
@@ 1 Introduction
Two topologies on regular operators.   The set $\mathcal{R}(H,H^{\prime})$ of regular operators has two important natural topologies. The Riesz topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
@@
from the norm topology on the space of bounded operators. Here $\bm{\chi}$ is the so called bounded transform map. The graph topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
$\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{P}(H\oplus H^{\prime})$
from the norm topology on the space of projections, where $\mathbf{p}$ is the map taking a regular operator to the orthogonal projection onto its graph. Let ${{}^{r}\mathcal{R}}(H,H^{\prime})$, resp. ${{}^{g}\mathcal{R}}(H,H^{\prime})$ denote the space of regular operators $H\to H^{\prime}$ equipped with the Riesz, resp. graph topology.
The Riesz topology is strictly stronger than the graph topology. In other words, the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is continuous, while the identity map ${{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ is not. On the subset $\mathcal{B}(H,H^{\prime})$ of bounded operators these two topologies coincide.
From graph to Riesz continuity.   Riesz continuous families of operators are much easier to deal with than mere graph continuous families. This motivates the approach that we are developing in this paper.
Suppose that one is given a graph continuous family of regular operators. One can try to modify it, in an appropriate way, so that the resulting family becomes Riesz continuous, then investigate the modified family and transfer results of this investigation to the original family. By an “appropriate way” we mean that such a transfer of results is possible, that is, some important properties of the original family may be recovered from the corresponding properties of the modified family. Of course, this description is quite vague; we will make it more precise below.
Let us illustrate our approach by an example of a family consisting of invertible operators. Recall that an operator $A\in\mathcal{R}(H)$ is called invertible if it is bijective as a map $\operatorname{dom}(A)\to H$. Such operator has a unique polar decomposition $A=u\cdot|A|$, where $u\in\mathcal{U}(H)$ is a unitary operator and $|A|=\sqrt{A^{*}A}$ is a positive operator. If $\mathcal{A}\colon X\to\mathcal{R}(H)$ is graph continuous, then the map $x\mapsto|\mathcal{A}_{x}|$ is Riesz continuous. If additionally all operators $\mathcal{A}_{x}$ are invertible, then the map $x\mapsto u_{x}=\mathcal{A}_{x}|\mathcal{A}_{x}|^{-1}$ from $X$ to ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is continuous. Taking $v_{x}=u_{x}^{*}$, we obtain a continuous map $v\colon X\to{{}^{s}{\mathcal{U}}}(H)$ such that the map $x\mapsto v_{x}\mathcal{A}_{x}=|\mathcal{A}_{x}|$ is Riesz continuous. The original family $\mathcal{A}$ can be recovered from the modified family $|\mathcal{A}|$ by the left multiplication by $u_{x}$.
This construction cannot be extended continuously to general non-invertible families because the polar decomposition of a general non-invertible operator involves a partial isometry which is not unitary. Moreover, this partial isometry does not depend continuously on the operator. However, one can still modify a family to a Riesz continuous one using the multiplication by an appropriate family of unitary operators, as we show in Theorem A below (we will multiply on the right, but it really does not matter: one can always pass to adjoint operators, replacing the left multiplication with the right one). In contrast with the construction above, the modified family no longer consists of positive operators, nor is it determined by the original family in any canonical way. But the modified family still carries an important information about the original one.
Theorem A.
There exists a map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map
@@
[…]
$K^{0}$ index for graph continuous families.   Let $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}$ be a graph continuous family and $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A. Then $\bm{\chi}\circ\Phi\circ\mathcal{A}=\varphi\circ\mathcal{A}$ is a norm continuous family of bounded operators. Suppose that all $\mathcal{A}_{x}$ are Fredholm. Then the operators $\varphi(\mathcal{A}_{x})$ are also Fredholm, and one can define the index of $\mathcal{A}$ as the index of this norm continuous family of bounded Fredholm operators:
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\varphi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X).$
This definition of the index is discussed in the last part of Section 3; it is justified by Theorem B below.
Equivalently, the index of $\mathcal{A}$ may be defined as the index of the family of projections $\Gamma_{x}\to H^{\prime}$ after a trivialization of the Hilbert bundle $\Gamma$ formed by the graphs $\Gamma_{x}$ of operators $\mathcal{A}_{x}$. (This trivialization should be chosen in the category of bundles with the structure group $\mathcal{U}$, not ${{}^{s}{\mathcal{U}}}$.) Note that such a trivialization is not canonical; it is defined only up to a continuous map $X\to\mathcal{U}(H)$. However, change of a trivialization does not affect the homotopy type of the corresponding family of bounded Fredholm operators, since the unitary group $\mathcal{U}(H)$ is contractible.
Searching for such a definition of the index was part of the motivation behind Theorem A. An alternative approach to the index of graph continuous families of Fredholm operators is provided by the previous paper [$P_{3}$] of the author. Another approach, for compact base spaces and in a more general framework of Hilbert modules over C*-algebras, is provided by results of M. Joachim [J]. Recently N. Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the index of families, which works under much weaker continuity assumptions than graph continuity. We show in the end of Section 3 that for graph continuous families of Fredholm operators our definition of the index given above coincides with Ivanov’s index.
Homotopy equivalences.   A regular operator $A$ is said to have compact resolvents***If the resolvent set $\operatorname{Res}(A)$ of $A$ is non-empty (for example, if $A$ is self-adjoint), then this definition agrees with the usual sense of the words “compact resolvent”: $(A-\lambda)^{-1}$ is compact for some (and then every) $\lambda\in\operatorname{Res}(A)$. However, a regular operator may have an empty resolvent set; the definition above covers such operators as well. if both $(1+A^{*}A)^{-1}$ and $(1+AA^{*})^{-1}$ are compact operators. Let $\mathcal{R}_{F}(H)$ and $\mathcal{R}_{K}(H)$ be the subsets of $\mathcal{R}(H)$ consisting of Fredholm operators and operators with compact resolvents, respectively. These subsets are invariant under the multiplication by unitary operators, so $\Phi(\mathcal{R}_{K})\subset\mathcal{R}_{K}$ and $\Phi(\mathcal{R}_{F})\subset\mathcal{R}_{F}$.
@@
are homotopy equivalences. See [$P_{3}$, Theorem A]. However, the proof in [$P_{3}$] provides no intuition on how a homotopy inverse map might look like. The following theorem reveals a possible form of homotopy inverse maps and provides an alternative proof of homotopy equivalence of the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
Theorem B.
@@
Let $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A. Then the restriction of $\Phi$ to the subspace of operators with compact resolvents is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$. Similarly, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$.
Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators. Then one can replace $\mathcal{R}_{K}$ in the statement of Theorem B with the set
$\mathcal{R}_{J}(H)=\left\{A\in\mathcal{R}(H)\,|\,(1+A^{*}A)^{-1},(1+AA^{*})^{-1}\in J\right\}.$
(In particular, $J$ can be the subset of trace class operators, or the subset of Hilbert–Schmidt operators.) Such a set $\mathcal{R}_{J}$ is invariant under the multiplication by unitary operators, so both $\Phi$ and homotopies used in the proof of Theorem B preserve $\mathcal{R}_{J}$. This provides the following generalization of the first part of Theorem B.
Theorem B’. Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators. Then the identity map ${{}^{r}\mathcal{R}}_{J}(H)\to{{}^{g}\mathcal{R}}_{J}(H)$ is a homotopy equivalence, with a homotopy inverse map given by $\Phi\colon{{}^{g}\mathcal{R}}_{J}\to{{}^{r}\mathcal{R}}_{J}$.
@@
Hilbert bundles.   Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be locally trivial Hilbert bundles over a topological space $X$ with a fiber $H$. The structure group of a general Hilbert bundle is the unitary group ${{}^{s}{\mathcal{U}}}(H)$ equipped with the strong operator topology. As was shown by Dixmier and Douady, every such bundle over a paracompact base space $X$ is trivial, that is, isomorphic to the trivial Hilbert bundle $X\times H\to X$. See [DD], Theorem 1. More generally, every numerable Hilbert bundle is trivial. However, automorphisms of the trivial Hilbert bundle $X\times H\to X$ are given by continuous maps $X\to{{}^{s}{\mathcal{U}}}(H)$ and preserve neither norm nor Riesz continuity of operator families, since the action of ${{}^{s}{\mathcal{U}}}(H)$ on $\mathcal{B}(H)$ by conjugation, $(u,a)\mapsto uau^{-1}$, is not continuous. Therefore, the notions of a norm continuous family of bounded operators or of a Riesz or graph continuous family of regular operators are not defined for a general Hilbert bundle.
@@
Nevertheless, the actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on the space $\mathcal{K}(H,H^{\prime})$ of compact operators are continuous, and thus the notion of a norm continuous family of compact operators $k_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ is well defined. By the same reason, the notion of a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents is well defined. Applying Theorem A to such a family, we obtain the following result.
Theorem C.
@@
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be numerable Hilbert bundles over a topological space $X$. Let $\mathcal{A}$ be a graph continuous family of regular operators $\mathcal{A}_{x}\in\mathcal{R}_{K}(\mathcal{H}_{x},\mathcal{H}^{\prime}_{x})$ with compact resolvents. Suppose that a global trivialization of one of these Hilbert bundles, $\mathcal{H}\cong X\times H$ or $\mathcal{H}^{\prime}\cong X\times H^{\prime}$, is fixed. Then there is a global trivialization of the second bundle taking $\mathcal{A}$ to a Riesz continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$.
We prove Theorems B and C in Section 3.
@@
Self-adjoint operators.   For self-adjoint operators, multiplication $A\mapsto Au$ or $A\mapsto uA$ should be replaced by the conjugation $A\mapsto uAu^{-1}$. Instead of independent trivializations of two Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, as in Theorem C, one should look for a one trivialization of a given Hilbert bundle $\mathcal{H}$. This narrows down available options drastically. In particular, Theorem A has no direct analogue for self-adjoint operators, even invertible ones. A graph continuous family of self-adjoint operators with compact resolvents may fail to be Riesz continuous in any trivialization of $\mathcal{H}$, as we will explain below. The next theorem shows that the obstruction to the existence of such a trivialization has purely local nature.
@@
From local to global trivializations.   Let $\mathcal{A}$ be a family of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ acting on fibers of a Hilbert bundle $\mathcal{H}$. Following Ivanov [$I_{2}$], we say that a trivialization $\tau\colon\left.\mathcal{H}\right|_{Y}\to Y\times H$ of $\mathcal{H}$ over a subset $Y\subset X$ is adapted to $\mathcal{A}$ if it takes $\mathcal{A}$ to a Riesz continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{R}(H)$. We say that $\mathcal{A}$ is Riesz quasi-continuous if, for every $x\in X$, it admits an adapted trivialization over a neighborhood of $x$.
Theorem D.
@@
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$, and let $\mathcal{A}$ be a family of self-adjoint operators with compact resolvents acting on fibers of $\mathcal{H}$. Suppose that $\mathcal{A}$ is Riesz quasi-continuous (that is, admits local adapted trivializations). Then there exists a global trivialization $\tau\colon\mathcal{H}\to X\times H$ adapted to $\mathcal{A}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
Applying the last statement of the theorem to the lifting of $\mathcal{A}$ to $X\times[0,1]$, we see that every two adapted trivializations are homotopic and thus the homotopy class of a Riesz continuous map $\tau_{*}\mathcal{A}\colon X\to\mathcal{R}_{K}^{\mathrm{sa}}$ does not depend on the choice of an adapted trivialization $\tau$. For a family $\mathcal{A}$ satisfying assumptions of the theorem, this allows to define the $K^{1}$ index of $\mathcal{A}$ as the index of the norm continuous family $\bm{\chi}(\tau_{*}\mathcal{A})$ of bounded Fredholm self-adjoint operators.
The main case of Theorem D is the case of families consisting of operators which are neither essentially positive nor essentially negative. In this case a similar trivialization theorem was recently proved by different methods and under much weaker continuity assumptions by N. Ivanov. See [$I_{2}$], Theorems 4.6 and 4.7. In combination with Lemma 5.1 from his paper [$I_{3}$], his trivialization theorem implies Theorem D for families consisting of operators which are neither essentially positive nor essentially negative.
Reading Ivanov’s paper [$I_{2}$] led the author to realizing that a part of the original proof of Theorem D can be separated from the rest and has an independent interest. In more details, Ivanov mentioned in the introduction to the first version of [$I_{2}$] that there is a notion of a polarization of a Hilbert bundle standing behind some proofs in [$I_{2}$] (he worked out the details in the current version of [$I_{2}$]). In Section 4 we introduce the notion of a polarization of a Hilbert bundle related to Ivanov’s polarizations in the same way as compact operators are related to finite rank ones, and prove an analogue of his trivialization theorem. See Theorem 4.1.
@@
From graph to Riesz continuity.   Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$, as above. The first obstruction to existence of an adapted trivialization is caused by the fact that $\mathcal{R}^{\mathrm{sa}}_{K}$ is path connected in the graph topology, but has three connected components in the Riesz topology. Two points from different components of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$ (for example, an essentially positive and an essentially negative operator) may be connected by a graph continuous path in $\mathcal{R}_{K}^{\mathrm{sa}}(H)$, but every trivialization of the trivial Hilbert bundle $[0,1]\times H$ will leave this path Riesz discontinuous.

### v3

@@ Abstract
We show that every graph continuous family of unbounded operators in a Hilbert space becomes Riesz continuous after one-sided multiplication by an appropriate family of unitary operators. This result provides a simple definition of the index for graph continuous families of Fredholm operators, and we show that for such families this index coincides with the index defined by N. Ivanov in [I2]. This result also has two corollaries for operators with compact resolvents: (1) the identity map between the space of such operators with the Riesz topology and the space of such operators with the graph topology is a homotopy equivalence; (2) every graph continuous family of such operators acting between fibers of Hilbert bundles becomes Riesz continuous in appropriate trivializations of the bundles.
For self-adjoint operators, multiplication by unitary operators should be replaced by conjugation. In general, a graph continuous family of self-adjoint operators with compact resolvents cannot be made Riesz continuous by an appropriate conjugation. We obtain a partial analogue of the “trivialization” result above for self-adjoint operators and describe obstructions to existence of such a trivialization in the general case. This motivates the notion of a polarization of a Hilbert bundle, and we prove a similar result for polarizations. These results are closely related to the recent work [I2] of Ivanov and provide alternative proofs for some of his results. We then show that, under a minor assumption on the space of parameters and for operators which are neither essentially positive nor essentially negative, there is always a trivialization making the family Riesz continuous.
footnotetext: Department of Mathematics, University of Haifa (Israel)
Department of Mathematics, Technion – Israel Institute of Technology
@@ 1 Introduction
Two topologies on regular operators.  The set $\mathcal{R}(H,H^{\prime})$ of regular operators has two important natural topologies. The Riesz topology on $\mathcal{R}(H,H^{\prime})$ is induced by the so called bounded transform map
@@
from the norm topology on the space of bounded operators. The graph topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
$\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{P}(H\oplus H^{\prime}),$
from the norm topology on the space $\mathcal{P}(H\oplus H^{\prime})$ of projections, where $\mathbf{p}$ is the map taking a regular operator to the orthogonal projection onto its graph. Let ${{}^{r}\mathcal{R}}(H,H^{\prime})$, resp. ${{}^{g}\mathcal{R}}(H,H^{\prime})$ denote the space of regular operators $H\to H^{\prime}$ equipped with the Riesz, resp. graph topology.
The Riesz topology is strictly finer than the graph topology. In other words, the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is continuous, while the identity map ${{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ is not. On the subset $\mathcal{B}(H,H^{\prime})$ of bounded operators these two topologies coincide.
From graph to Riesz continuity.  Using the bounded transform, one can deal with Riesz continuous families of operators in essentially the same manner as with norm continuous families of bounded operators. Because of that, Riesz continuous families are much easier to deal with than mere graph continuous families. Our first result shows that one can turn a graph continuous family to a Riesz continuous one by a “change of coordinates”***The interpretation of our construction as a change of coordinates was suggested by the referee..
Theorem 1.
There exist a map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map
@@
Passing to adjoint operators, one obtains a similar statement for the left multiplication. Namely, there is a map $v^{\prime}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H^{\prime})$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map $A\mapsto v^{\prime}(A)\cdot A$ is graph-to-Riesz continuous.
We will now present three applications of Theorem 1.
$K^{0}$ index for graph continuous families.  Let $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}$ be a graph continuous map and $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem 1. Then the composition
$\alpha=\bm{\chi}\circ\Phi\circ\mathcal{A}\colon X\to\mathcal{B}(H)$
is norm continuous. Suppose that all $\mathcal{A}_{x}$ are Fredholm. Then the operators $\alpha_{x}$ are also Fredholm, and one can define the index of $\mathcal{A}$ as the classical index of $\alpha\colon X\to\mathcal{B}_{F}$ (here $\mathcal{B}_{F}$ denotes the space of bounded Fredholm operators):
$\operatorname{ind}\mathcal{A}=\operatorname{ind}\alpha=[\alpha]\in[X,\mathcal{B}_{F}]=K^{0}(X).$
This definition of the index is justified by Theorem 2 below. We discuss the index of graph continuous families in the last part of Section 4.
Searching for such a definition of the index was part of the motivation behind Theorem 1. An alternative approach to the index of graph continuous families of Fredholm operators is provided by the previous paper [P3] of the author. Another approach, for compact base spaces and in a more general framework of Hilbert modules over C*-algebras, is provided by results of M. Joachim [J]. Recently N. Ivanov developed a new approach to the index of families [I1, I2], which works under much weaker continuity assumptions than graph continuity. We show in the end of Section 4 that the definition of the index given above coincides with Ivanov’s index for graph continuous families of Fredholm operators.
Homotopy equivalences.  A regular operator $A$ is said to have compact resolventsIf the resolvent set $\operatorname{Res}(A)$ of $A$ is non-empty (for example, if $A$ is self-adjoint), then this definition agrees with the usual sense of the words “compact resolvent”: $(A-\lambda)^{-1}$ is compact for some (and then every) $\lambda\in\operatorname{Res}(A)$. However, a regular operator may have an empty resolvent set; the definition above covers such operators as well. if both $(1+A^{*}A)^{-1}$ and $(1+AA^{*})^{-1}$ are compact operators. Let $\mathcal{R}_{F}(H)$ and $\mathcal{R}_{K}(H)$ be the subsets of $\mathcal{R}(H)$ consisting of Fredholm operators and operators with compact resolvents, respectively. These subsets are invariant under the multiplication by unitary operators, so $\Phi(\mathcal{R}_{K})\subset\mathcal{R}_{K}$ and $\Phi(\mathcal{R}_{F})\subset\mathcal{R}_{F}$.
@@
are homotopy equivalences. See [P3, Theorem A]. However, the proof in [P3] provides no intuition on how a homotopy inverse map might look like. The following theorem reveals a possible form of homotopy inverse maps and provides an alternative proof of homotopy equivalence of the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
Theorem 2.
@@
Let $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem 1. Then the restriction of $\Phi$ to the subspace of operators with compact resolvents is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$. Similarly, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$.
@@
Hilbert bundles.  Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be locally trivial Hilbert bundles over a topological space $X$ with a fiber $H$. (In the following we will always suppose that all Hilbert bundles are locally trivial.) The structure group of a general Hilbert bundle is ${{}^{s}{\mathcal{U}}}(H)$, the unitary group of $H$ with the strong operator topology. This group is contractible [DD, Lemma 3], so every such bundle over a paracompact base space $X$ is trivial, that is, isomorphic to the trivial Hilbert bundle $X\times H\to X$. More generally, every numerable Hilbert bundle is trivial. However, automorphisms of the trivial Hilbert bundle $X\times H\to X$ are given by continuous maps $X\to{{}^{s}{\mathcal{U}}}(H)$ and preserve neither norm nor Riesz continuity of operator families, since the action of ${{}^{s}{\mathcal{U}}}(H)$ on $\mathcal{B}(H)$ by conjugation, $(u,a)\mapsto uau^{-1}$, is not continuous. Therefore, the notions of a norm continuous family of bounded operators or a Riesz/graph continuous family of regular operators are not defined for a general Hilbert bundle.
@@
Nevertheless, the actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on the space $\mathcal{K}(H,H^{\prime})$ of compact operators are continuous, and thus the notion of a norm continuous family of compact operators $k_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ is well defined. By the same reason, the notion of a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents is well defined (Proposition 2.3). Applying Theorem 1 to such a family, we obtain the following result.
Theorem 3.
@@
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be numerable Hilbert bundles over a topological space $X$. Let $\mathcal{A}$ be a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents. Suppose that a trivialization of one of these Hilbert bundles, $\mathcal{H}\cong X\times H$ or $\mathcal{H}^{\prime}\cong X\times H^{\prime}$, is fixed. Then there is a trivialization of the second bundle taking $\mathcal{A}$ to a Riesz continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$.
@@
Self-adjoint case.  For self-adjoint operators, multiplication $A\mapsto Au$ or $A\mapsto uA$ should be replaced by the conjugation $A\mapsto uAu^{-1}$. Instead of independent trivializations of two Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, as in Theorem 3, one should look for a one trivialization of a given Hilbert bundle $\mathcal{H}$. This narrows down available options drastically. In particular, Theorem 1 has no direct analogue for self-adjoint operators, even invertible ones.
@@
Let $\mathcal{A}$ be a family of regular operators acting on fibers of a locally trivial Hilbert bundle $\mathcal{H}$ over $X$. Following Ivanov [I3], we say that a trivialization $\tau\colon\left.\mathcal{H}\right|_{Y}\to Y\times H$ of $\mathcal{H}$ over a subset $Y\subset X$ is adapted“Fully adapted” in terms of Ivanov to $\mathcal{A}$ if it takes $\mathcal{A}$ to a Riesz continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{R}(H)$. In particular, $\tau$ is adapted to a family $\mathcal{A}$ of bounded operators if it takes $\mathcal{A}$ to a norm continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{B}(H)$.
Spectral projections of a graph continuous family $\mathcal{A}$ of self-adjoint operators with compact resolvents give rise to what we call a compatible family of local bundle projections. A trivialization is adapted to $\mathcal{A}$ if and only if it is adapted to these projections. Our main results in the self-adjoint case, Theorems 4 and 5, are concerned with such families of projections, and we deduce results concerning self-adjoint operators from these two theorems.
Polarizations.  We say that $p\colon\left.\mathcal{H}\right|_{Y}\to\left.\mathcal{H}\right|_{Y}$ is a bundle projection of $\mathcal{H}$ over $Y\subset X$ if $p$ is a continuous bundle map such that each $p_{x}$ is a projection. We say that two bundle projections, $p$ over $Y$ and $p^{\prime}$ over $Y^{\prime}$, are compatible (resp. finitely compatible) if the difference $x\mapsto p_{x}-p^{\prime}_{x}$ is a norm continuous family of compact (resp. finite rank) operators over $Y\cap Y^{\prime}$.
Locally, a polarization is defined by a bundle projection $p$ over an open subset $Y\subset X$ such that each $p_{x}$ has infinite-dimensional range and kernel. Globally, a polarization is defined by an open covering $(X_{i})$ of $X$ together with a collection of pairwise compatible local polarizations $p_{i}$ over $X_{i}$. Following Ivanov [I2], we call such a collection $(X_{i},p_{i})$ a polarization atlas and define a polarization as an equivalence class of polarization atlases (as usual, two atlases are equivalent if their union is also an atlas). A finite polarization is defined in the same way, but with “compatible” replaced by “finitely compatible”. See Section 5 for details.
Let $\Pi$ be a polarization of $\mathcal{H}$ given by an atlas $(X_{i},p_{i})$. We say that a (local or global) trivialization of $\mathcal{H}$ is adapted to $\Pi$ if it is adapted to each $p_{i}$ (that is, takes each $p_{i}$ to a norm continuous family). We say that a polarization is tame if it admits a local adapted trivialization over a neighborhood of each point $x\in X$.
The notion of a polarization introduced by Ivanov in the appendix to [I2] is essentially equivalent to a tame finite polarization in our terminology.
Theorem 4.
The next theorem shows that a modest assumption on the base space is enough for existing of an adapted trivialization.
Let us say that a paracompact space $X$ has locally finite covering dimension if each point of $X$ has a neighborhood of finite covering dimension (we do not require this neighborhood to be open or small). In particular, every space of finite covering dimension has locally finite covering dimension.
Theorem 5.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ of locally finite covering dimension. Then every polarization of $\mathcal{H}$ is tame and thus admits a global adapted trivialization.
Self-adjoint operators and adapted trivializations.  We want to know whether a family $\mathcal{A}$ of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$ admits a global adapted trivialization. Clearly, for that to be true, $\mathcal{A}$ has to admit local adapted trivializations (by that we mean that for every $x\in X$ there is an adapted trivialization over a neighborhood of $x$). Theorem 4 implies the following result, which shows that for essentially unitary operators this local condition is also sufficient. The author does not know whether this is true for arbitrary bounded self-adjoint operators, even if they are Fredholm.
Theorem 6.
@@
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$, $\mathcal{A}$ be a family of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$, and either all operators $\mathcal{A}_{x}$ are essentially unitary or they all have compact resolvents. Suppose that $\mathcal{A}$ admits local adapted trivializations. Then $\mathcal{A}$ admits a global adapted trivialization. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization. In particular, every two such trivializations are homotopic (in the space of trivializations adapted to $\mathcal{A}$).
The main case of Theorem 6 is the case of families consisting of operators which are neither essentially positive nor essentially negative. In this case, a similar trivialization theorem was nearly simultaneously proved by different methods and under much weaker continuity assumptions by Ivanov. See [I2], Theorems 4.5–4.7. It can be easily shown that his Theorems 4.6–4.7 imply Theorem 6 for families consisting of operators which are neither essentially positive nor essentially negative. Ivanov mentioned in the introduction to the first version of [I2] that there is a notion of a polarization of a Hilbert bundle standing behind some proofs in [I2]. This led the author to introducing her version of a polarization of a Hilbert bundle and proving an analogue of Ivanov’s trivialization theorems 4.6–4.7 in terms of polarizations (Theorem 4). Later, Ivanov wrote down the details of his notion of a polarization in the appendix to the current version of [I2] and rephrased the proofs of his Theorems 4.5–4.7 in terms of polarizations. See [I2], Theorems A.2–A.4.
@@
First local obstruction.  Let $\mathcal{R}^{\mathrm{sa}}_{K}(H)$ denote the set of self-adjoint operators with compact resolvents. The first local obstruction to existence of an adapted trivialization for a graph continuous family $\mathcal{A}$ of operators $\mathcal{A}_{x}\in\mathcal{R}^{\mathrm{sa}}_{K}(\mathcal{H}_{x})$ is caused by the fact that $\mathcal{R}^{\mathrm{sa}}_{K}$ is path connected in the graph topology, but has three connected components in the Riesz topology. Two points from different components of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$ (for example, an essentially positive and an essentially negative operator) may be connected by a graph continuous path in $\mathcal{R}_{K}^{\mathrm{sa}}(H)$, but no trivialization of the trivial Hilbert bundle $0,1]\times H$ makes this path Riesz continuous.